# LakeVilla: Multi-Table Transactions for Lakehouses

### v1

@@ 7. Conclusion
Current lakehouses provide limited transaction support without auxiliary structures. In this work, we introduce LakeVilla to improve transactional models and guarantees of state-of-the-art lakehouses. LakeVilla defines three modular features to achieve its goal and provides transactional guarantees up to serializability. Our results show the trade-offs and compatibility with systems used in practice. In contrast to other solutions in the Lakehouse area, we extend the base open table format on the object store non-invasively and allow transactions to freely choose their required guarantees, opening up interesting directions for future architectures and approaches.
@@ Acknowledgements.
We are grateful to the colleagues at the Chair of Database Systems at TUM, the SAP HANA Cloud team, and all members of the SAP HANA Campus for their feedback. This work was partly funded by the Chair of Database Systems at TUM and SAP.

### v2

@@ 8 Conclusion
Current lakehouses provide limited transaction support without auxiliary structures. In this work, we introduce LakeVilla to improve transactional models and guarantees of state-of-the-art lakehouses. LakeVilla defines three modular features to achieve its goal and provides transactional guarantees up to serializability. Our results show the trade-offs and compatibility with systems used in practice. In contrast to other solutions in the Lakehouse area, we extend the base open table format on the object store non-invasively and allow transactions to freely choose their guarantees, opening up interesting directions for future architectures and approaches.
9 Statements and Declarations
Funding. This work was funded by the Chair of Database Systems at TUM and SAP SE. The PhD position of Tobias Götz is partially funded by SAP SE.
@@ Acknowledgements.
We are grateful to the colleagues at the Chair of Database Systems at TUM, the SAP HANA Cloud team, and all members of the SAP HANA Campus for their feedback.